If 42 ounces of red beads that sell for $7 per ounce are mixed with 58 ounces of golden beads that sell for $10 per ounce, what should be the selling price for the mixture of beads?
Answer by Maths68(1474)


Red Beads = 42 oz
Price = $7/oz
Golden Beads = 58 oz
Price = 10/oz
Mixture = (42+58)oz=100oz
Pirce = x = $?/oz


(Mixture)(Price of mixture)=(Red Beads)(Price of Red Beads)+(Golden Beads)(Price of Golden Beads)
(100)x=(42*7)+(58*10)
100x=294+580
100x=874
x=874/100
x=8.74


$8.74 per ounce should be the selling price for the mixture of beads.